# VoxCity: A Seamless Framework for Open Geospatial Data Integration, Grid-Based Semantic 3D City Model Generation, and Urban Environment Simulation

### v1

@@ 3.3 Tree canopy height
We screened tree canopy height datasets according to: (1) global coverage, (2) resolution finer than 10 m, (3) coverage that includes urban areas, and (4) no restriction to specific tree species. During this screening, we excluded datasets that cover only particular cities or nations [122], have resolutions coarser than 10 m [123], or focus on forested regions or specific tree species [124, 125]. We also considered open tree inventories, which typically include a database of individual trees (e.g., location, size, age, species) [126, 127, 128]. However, each city or country tends to maintain its own tree inventory format, and coverage rarely extends beyond one locality [129, 130, 131]. Hence, no standardized globally comprehensive tree inventory dataset was identified. Ultimately, we selected High Resolution 1 m Global Canopy Height Maps (META) [23] and ETH Global Sentinel-2 10 m Canopy Height (2020) (ETH) [132] for further review. Table 4 provides details on these data sources.
@@ 3.4 Terrain elevation
We filtered Digital Elevation Model (DEM) and Digital Terrain Model (DTM) sources according to: (1) a horizontal resolution of at least 1 Arc second (30.9 m at the equator), and (2) a bare-earth model that excludes buildings and vegetation (i.e., a DTM rather than a Digital Surface Model, DSM). Datasets with 3 Arc seconds resolution [133], as well as DEMs or DSMs that retain building and vegetation heights [134, 135, 136, 137] were excluded. Consequently, we selected FABDEM [25], DeltaDTM [138], USGS 3DEP 1m DEM (USDEM) [139], England 1m Composite DTM (ENGDTM) [140], Australian 5M DEM (AUSDEM) [141], and RGE Alti (FRADEM) [142] for further review.
@@ 3.5 Data catalog with guidelines for data selection
(1) Use EUBUCCO for EU countries; for other regions, employ OSM or Overtures as the base source.
(2) If the target area is covered by any complementary dataset (MSBF for the USA, Europe, and Australia; OB2.5DT for the Global South; UT-GLOBUS for 1,200 global cities), use it as the complementary source.
@@
Terrain elevation USDEM [139], ENGDTM [140], AUSDEM [141], DeltaDTM [138], FABDEM [25] (1) For cities covered by high-resolution datasets (ENGDTM, USDEM, AUSDEM, and FRADEM), use those data sources.
@@ 4 The VoxCity package
In this section, we introduce ‘VoxCity’, an open-source Python package for open geospatial data integration, grid-based 3D city model generation, and urban environment simulation. Figure 6 illustrates the framework of VoxCity. Users start the process by specifying the target area and voxel size and selecting data sources from our catalog. The workflow consists of four main sub-processes: (1) 3D city model generation, (2) simulation, (3) file export, and (4) visualization. In (1) 3D city model generation, VoxCity downloads building height, land cover, tree canopy height, and terrain elevation data from the selected sources within the specified target area. It then voxelizes all the downloaded data and integrates them into a semantic 3D city model. In (2) simulation, the output models can be used directly to conduct urban environment simulations through VoxCity’s built-in simulation functions. In (3) file export, VoxCity can export the output model in multiple file formats that are compatible with external software. In (4) visualization, VoxCity offers 3D visualization functionality for not only generated city models but also simulation results. The following subsections provide more details on these sub-processes.
@@ 4.1.1 Download
VoxCity automatically downloads the required data within a target area through either file links or API depending on selected data sources. Users can define the target area as a rectangular region. Downloaded data are saved in users’ environments and utilized in subsequent processes. Vector and raster data are saved as GeoJSON and GeoTIFF files, respectively. Users who need these intermediate files can also use VoxCity solely as a downloader for various geospatial datasets, similar to OSMnx [96], which often serves as a downloader for road networks from OSM.
@@ 4.1.2 Voxelization and integration
VoxCity first aggregates values using a two-dimensional horizontal grid defined by the voxel size (in meters). For building height, tree canopy height, and land cover, each cell’s value is determined by the dominant value within that cell. For instance, focusing on land cover, when a cell includes multiple classes, the class that covers the largest area is assigned as its representative value of the cell. In this process, the land cover class harmonization shown in Table 3 is applied, enabling users to easily compare the generated 3D models across different data sources and select the desirable sources. For terrain elevation, the representative value of a cell is calculated as the average within the cell.
@@
VoxCity’s 3D city models are structured as three-dimensional arrays, whose dimensions correspond to the geospatial x, y, and z axes. Each cell corresponds to a voxel, and its value represents an element type such as a building, tree, or water.
@@ 4.1.3 Example outputs
Additionally, VoxCity’s outputs capture the morphological characteristics of each city by combining building shapes and arrangements, land covers, tree canopies, and terrain elevations. Notable examples include the radiating block arrangements and street networks around Arc de Triomphe in Paris; dense skyscrapers surrounded by water in Lower Manhattan, New York City; tightly packed houses on a mountainside in Rio de Janeiro; and sparse high-rise buildings set amid sandy terrain in Abu Dhabi.
The approach is efficient. In our environment—equipped with an Intel Core i9-13900 processor and an internet connection with a measured download speed of 180 Mbps—VoxCity took 25 seconds to download the required data and generate the 3D city model of Singapore shown in Figure 7. However, it is important to note that processing times can vary substantially depending on the selected data sources.
@@ 4.2.1 Solar radiation
A module, ‘solar’, provides functionality for calculating global irradiance at ground-level and on building surfaces. Solar radiation is a critical factor influencing outdoor heat stress [143], thermal comfort [144, 145], urban farming [146], photovoltaic power generation [147, 148], and building energy consumption by HVAC and lighting [149, 150]. Consequently, it has profound impacts on urban sustainability and the well-being of city dwellers—underscoring its importance in urban and architectural planning. The ‘solar’ module offers two calculation options: (1) instantaneous solar irradiance (Wm-2) at a specific timestamp and (2) cumulative solar irradiance (Whm-2) over a specific period. For instantaneous value, the module first calculates solar azimuth and elevation angles using the location (longitude and latitude) of the target area and the target timestamp, employing a Python package ‘astral’ [^12] https://github.com/sffjunkie/astral. Direct and diffuse irradiance are then calculated using a ray-tracing technique, incorporating the computed sun position as well as direct normal and diffuse horizontal irradiance that is not affected by shading. In the ray-tracing process, a transmittance determined by the Beer-Lambert law (Equation 1) is incorporated when rays pass through tree voxels:
[…]
For cumulative value, the module calculates instantaneous irradiance at each timestamp from the start to the end of a specified period and then sums the results. To incorporate locally relevant climate conditions (including direct normal and diffuse horizontal irradiance), the module can use Energy Plus Weather (EPW) files. Users may supply their own EPW files; otherwise, the module can automatically download the nearest EPW file from Climate.OneBuilding.Org  [^13] https://climate.onebuilding.org/.
@@
Additionally, VoxCity provides a feature to aggregate grid-based simulation results along the edges of road networks, as shown in Figure 8b. The functionality uses OSMnx to download road networks from OSM, enabling various network analyses within simulated urban environments—including not only solar irradiance but also view indices and landmark visibility, which is detailed in Section 4.2.2.
@@ 4.2.2 View index
A module, ‘view’, provides functionality for conducting view index analyses, where the ratio of specific object types—such as vegetation, sky, and buildings—visible from a given location is quantitatively evaluated. The module calculates the view index by applying a ray-tracing technique to 3D city models. Users can specify the target object type, the vertical angle of view (i.e., the ranges in which rays are cast), and the total number of rays. The horizontal angle of view is set to 360 degrees, assuming uniform visibility in all directions.
@@
The view module also provides functionality for evaluating the visibility of specified landmark buildings from ground-level locations using a ray-tracing technique. This enables users to understand from which locations within a target area the landmark buildings can be seen. Figure 8e shows example visibility maps. Landmarks play important roles in pedestrians’ perception and wayfinding [152, 153], influencing visual comfort and walkability [154, 155]. The access to the view of landmarks is proved to adding property value [156]. The visibility simulations could support informed urban planning considering such effects of landmarks.
@@ 4.3 File export
VoxCity’s ‘exporter’ subpackage can save the generated 3D city model data in several file formats to support downstream applications and data exchange. The available file formats include INX for ENVI-met [157, 158], OBJ for 3D modeling software, VOX for MagicaVoxel [^14] https://ephtracy.github.io/index.html?page=mv_main, and NPY for Python. ENVI-met is a widely used microclimate simulation software. MagicaVoxel offers GPU-accelerated 3D rendering for voxel models. NPY files enable model data transfer between different Python environments without any format conversion. When using the NPY format, the data is directly saved as a three-dimensional NumPy [159] array. For other formats, the data is first converted to each format’s specific data structure. In the case of VOX, the 3D array from VoxCity is reformatted into the VOX structure, including color palette information corresponding to voxel classes. The subsequent subsections describe the export functionality for the INX and OBJ formats in more detail.
@@ 4.3.1 INX for ENVI-met
Building height and terrain elevation from VoxCity are directly converted to the INX format, while canopy height and land cover data are translated into corresponding vegetation ID and material ID in ENVI-met before being integrated into the INX format. Additionally, users can specify the trunk height ratio relative to total canopy height and the leaf area density (LAD) of trees. These tree-specific settings are then exported separately as a project database (EDB) file.
@@
Figure 9 shows microclimate simulation examples using the widely-used ENVI-met code (V5.7.1, ENVI-met Gmbh, Essen) in two characteristic sites, high-rise Al Zahiya district in Abu Dhabi and low-rise residential in Greenbelt, MD (USA). Since ENVI-met is based on the structured mesh (i.e., Cartesian coordinate), VoxCity can provide basic 3D geometries (Figures 9a and 9b) for the simulation. The calculation domains were 800 m × 800 m × 310 m for Abu Dhabi, and 500 m × 500 m × 50 m for Greenbelt. They comprised inner areas with buildings and trees with horizontally homogeneous grid spacing (4 m for Abu Dhabi and 2 m for Greenbelt here), which is off-the-shelf from VoxCity. Because perimeter margins are necessary for ENVI-met simulation to avoid unphysical results (as detailed in their documentation: https://envi-met.com/tutorials-plugins-faqs-helpful-info/), VoxCity provides a function to remove perimeter buildings and trees. For vertical grid spacing, VoxCity provides options that use the same size for the horizontal grid or set an arbitrary number of layers. Here, we set 19 layers for the Greenbelt case and 50 for the Abu Dhabi case. One should note that before running the simulation, ENVI-met automatically resizes the meshes within the five layers near the ground to a finer resolution. We also note that ENVI-met V5.7.1 often predicts unphysical results for complex terrains (as detailed on their website: http://www.envi-hq.com/), so we strongly recommend VoxCity users not to obtain DEM when conducting ENVI-met simulation.
As mentioned above, VoxCity automatically sets default values of the thermal properties of buildings, ground, and trees by providing the INX file. Although overwriting the INX file, e.g., setting detailed building surface materials, can give more accurate predictions, we show the results with the default values here to verify their adequacy. For meteorological forcings, we used the 2004–2018 typical meteorological year data of the Energy Plus Weather file (https://www.ladybug.tools/epwmap/) and selected the nearest weather stations for each site.
Figures 9c–9f show thermal environments at 1.5 m height on typical summer afternoons (2:00 p.m.), namely the horizontal wind vectors and the universal thermal climate index (UTCI) (derived from the air temperature, humidity, radiation, and wind speed) predicted by ENVI-met. In the Abu Dhabi case, the inlet wind was the northwesterly sea breeze from the Persian Gulf. The speed was relatively high above the roads and the open areas and reached 5 m/s at the maximum (Figure 9c), probably because the city block pattern is along the direction of the inlet wind. However, the inlet air temperature was 44°C, hotter than the human body, and therefore, as seen around point ① in Figure 9e for example, the increase in wind speed rather increased UTCI by convective heat transfer (see Fig. 7 of Bröde et al. [160]). Shading effects of buildings and trees on UTCI are evident, and the difference between sunlit and shaded areas reached $>$ 20 °C (Figure 9e). Weak and complex diverged flows with lower air temperature and UTCI values are found in the leewards of the buildings (e.g., around building ② in Figures 9c and  9e and the southeast city block), indicating cooler above-roof air moved to the ground level by the downdrafts. In the Greenbelt case, the wind speed was generally lower than the inlet value (1.5 m/s) due to the momentum sink by trees (Figure 9d). The UTCI distribution seems to correspond simply to the shading pattern, and the contrast between roads and under-tree areas is evident (Figure 9f). Overall, these results seem to be reasonable and suggest the adequacy of the INX file obtained from VoxCity.

### v3

@@ 3.3 Tree canopy height
To screen datasets, we set criteria: (1) coverage of urban areas and (2) no restriction to specific tree species, in addition to the aforementioned criteria for spatial coverage and resolution. Therefore, we excluded datasets that focus on forested regions or specific tree species [116, 117]. We also considered open tree inventories, which typically include a database of individual trees (e.g., location, size, age, species) [118, 119, 120]. However, each city or country tends to maintain its own tree inventory format, and coverage rarely extends beyond one locality [121, 122, 123]. Hence, no standardized globally comprehensive tree inventory dataset was identified. Ultimately, we selected the High Resolution 1 m Global Canopy Height Maps (META) [23] and ETH Global Sentinel-2 10 m Canopy Height (2020) (ETH) [124] for further review. Table 4 provides details on these data sources.
@@ 3.4 Terrain elevation
To select terrain elevation datasets, we set a resolution threshold of 1 Arc second (30.9 m at the equator) instead of 10 m for other data types. This exception was made because, for terrain elevation data with global coverage, only datasets with resolutions coarser than 1 Arc second are available. We allowed this coarser resolution because terrain elevation in urban areas rarely changes drastically at spatial scales less than 30 m. Additionally, we established a criterion requiring a bare-earth model to exclude DEMs or DSMs that retain building and vegetation heights [125, 126, 127, 128]. Consequently, we selected Forest And Buildings removed Copernicus 30m DEM (FABDEM) [25], DeltaDTM [129], USGS 3DEP 1m DEM (USDEM) [130], England 1m Composite DTM (ENGDTM) [131], Australian 5M DEM (AUSDEM) [132], and RGE Alti (FRADEM) [133] for further review. While USDEM, ENGDTM, AUSDEM, and FRADEM have only national coverage, they were incorporated because of their exceptionally fine spatial resolution, ranging from 1 to 5 m.
@@ 3.5 Data catalog with guidelines for data selection
(1) Use EUBUCCO for EU countries; for other regions, employ OSM or Overture as the base source.
(2) If the target area is covered by any complementary dataset (MSBF for the USA, Europe, and Australia; OB2.5DT for Africa, South Asia, South-East Asia, Latin America and the Caribbean; UT-GLOBUS for 1,200 global cities), use it as the complementary source.
@@
Terrain elevation USDEM [130], ENGDTM [131], AUSDEM [132], FRADEM [133], DeltaDTM [129], FABDEM [25] (1) For cities covered by high-resolution datasets (ENGDTM, USDEM, AUSDEM, and FRADEM), use those data sources.
@@ 4 The VoxCity package
In this section, we introduce ‘VoxCity’, an open-source Python package for open geospatial data integration, grid-based 3D city model generation, and urban environment simulation. Figure 6 illustrates the framework of VoxCity. Users start the process by specifying the target area and voxel size and selecting data sources from our catalog. The workflow consists of four main sub-processes: (1) 3D city model generation, (2) simulation, (3) file export, and (4) visualization. In (1) 3D city model generation, VoxCity downloads building height, land cover, tree canopy height, and terrain elevation data from the selected sources within the specified target area. It then voxelizes all the downloaded data and integrates them into a semantic 3D city model. In (2) simulation, the output models can be used directly to conduct urban environment simulations through VoxCity’s built-in simulation functions. In (3) file export, VoxCity can export the output model in multiple file formats that are compatible with external software. In (4) visualization, VoxCity offers 3D visualization functionality for not only generated city models but also simulation results.
@@ 4.1.1 Download
VoxCity automatically downloads the required data within a target area through either file links or APIs depending on selected data sources. Users can define the target area as a rectangular region. Downloaded data are saved in users’ environments and utilized in subsequent processes. Vector and raster data are saved as GeoJSON and GeoTIFF files, respectively. Users who need these intermediate files can also use VoxCity solely as a downloader for various geospatial datasets, similar to OSMnx [99], which often serves as a downloader for road networks from OSM.
@@ 4.1.2 Voxelization and integration
VoxCity first aggregates values using a two-dimensional horizontal grid defined by the voxel size (in meters). For building height, tree canopy height, and land cover, each cell’s value is determined by the dominant value within that cell. For instance, focusing on land cover, when a cell includes multiple classes, the class that covers the largest area is assigned as the representative value of the cell. In this process, the land cover class harmonization shown in Table 3 is applied, enabling users to easily compare the generated 3D models across different data sources and select the desirable sources. For terrain elevation, the representative value of a cell is calculated as the average within the cell.
@@
VoxCity’s 3D city models are structured as three-dimensional arrays, whose dimensions correspond to the geospatial x, y, and z axes as shown in Figure 13 in Appendix. Each cell corresponds to a voxel, and its value represents an element type such as a building, tree, or water. While some previous studies implemented octrees for voxel model structures to enhance efficiency in data processing [36, 134], VoxCity does not currently use this approach. Incorporating octrees could reduce memory usage and improve ray-tracing-based simulation performance; therefore, we plan to implement this structure in future work.
@@ 4.1.3 Example outputs
VoxCity’s outputs capture the morphological characteristics of each city by combining building shapes and arrangements, land cover, tree canopies, and terrain elevations. Notable examples include the radiating block arrangements and street networks around Arc de Triomphe in Paris; dense skyscrapers surrounded by water in Lower Manhattan, New York City; tightly packed houses on a mountainside in Rio de Janeiro; and sparse high-rise buildings set amid sandy terrain in Abu Dhabi. Additionally, to further demonstrate VoxCity’s capability for diverse geographies worldwide, we generated models for eight cities: Vancouver, San Francisco, Medellín, La Paz, Barcelona, Nairobi, Mumbai, and Bangkok (Figure 14 in Appendix).
To demonstrate the efficiency of our approach, we evaluated the time requirements for manual and computational processing to generate 3D city models, as shown in Table 8 in Appendix. Manual processing includes specifying the target area and setting data sources, while computational processing includes downloading the required data and generating the voxel city models. In our test environment—equipped with an Intel Core i9-13900 processor and an internet connection with a measured download speed of 180 Mbps—VoxCity required 90 seconds for manual and 40 seconds for computational processing for the New York City model shown in Figure 7. Computational processing times can vary substantially depending on the urban morphology and selected data sources. For instance, generation of the Singapore model shown in Figure 7, which exhibits lower building density than that of New York City, required a computational processing time of 25 seconds.
@@ 4.2.1 Solar radiation
A module, ‘solar’, provides functionality for calculating global irradiance at ground-level and on building surfaces. Solar radiation is a critical factor influencing outdoor heat stress [135], thermal comfort [136, 137], urban farming [138], photovoltaic power generation [139, 140], and building energy consumption by HVAC and lighting [141, 142]. Consequently, it has profound impacts on urban sustainability and the well-being of city dwellers—underscoring its importance in urban and architectural planning. We employ a ray-tracing based calculation method for solar irradiance introduced by Pružinec and Ďuračiová [143], Monsi and Saeki [144]. The method is summarized in D. The ‘solar’ module offers two calculation options: (1) instantaneous solar irradiance (Wm-2) at a specific timestamp and (2) cumulative solar irradiance (Whm-2) over a specific period, such as a day, a month, or a year.
@@
Additionally, VoxCity provides a feature to aggregate grid-based simulation results along the edges of road networks, as shown in Figure 8b. The functionality uses OSMnx to download road networks from OSM, enabling various network analyses within simulated urban environments—including not only solar irradiance but also view indices and landmark visibility, which are detailed in Section 4.2.2.
@@ 4.2.2 View index
A module, ‘view’, provides functionality for conducting view index analyses, where the ratio of specific object types—such as vegetation, sky, and buildings—visible from a given location is quantitatively evaluated. The module calculates the view index by applying a ray-tracing technique to 3D city models. Specifically, we implemented a line of sight evaluation method introduced by Mor et al. [145] for voxel city models. Users can specify the target object type, the vertical angle of view (i.e., the ranges in which rays are cast), and the total number of rays. The horizontal angle of view is set to 360 degrees, assuming uniform visibility in all directions.
@@
The view module also provides functionality for evaluating the visibility of specified landmark buildings from ground-level locations using a ray-tracing technique. Specifically, we implemented the ground-level viewshed calculation introduced by Wróżyński et al. [146] for voxel city models. This enables users to understand from which locations within a target area the landmark buildings can be seen. Figure 8e shows example visibility maps. Landmarks play important roles in pedestrians’ perception and wayfinding [147, 148], influencing visual comfort and walkability [149, 150]. The access to the view of landmarks has been proved to add property value [151]. The visibility simulations could support informed urban planning considering such effects of landmarks.
@@ 4.3 File export
VoxCity’s ‘exporter’ subpackage can save the generated 3D city model data in several file formats to support downstream applications and data exchange. The available file formats include INX for ENVI-met [152, 153], the Wavefront OBJ for 3D modeling software, the MagicaVoxel [^13] https://ephtracy.github.io/index.html voxel model format (VOX), and the Pickle (PKL) for Python. ENVI-met is a widely used microclimate simulation software. MagicaVoxel offers GPU-accelerated 3D rendering for voxel models. PKL files enable model data transfer between different Python environments without any format conversion. When using the PKL format, the data is directly saved as a three-dimensional NumPy [154] array. For other formats, the data is first converted to each format’s specific data structure. In the case of VOX, the 3D array from VoxCity is reformatted into the VOX structure, including color palette information corresponding to voxel classes. The subsequent subsections describe the export functionality for the INX and OBJ formats in more detail.
@@ 4.3.1 INX for ENVI-met
Building height and terrain elevation from VoxCity are directly converted to the INX format, while canopy height and land cover data are translated into corresponding vegetation ID and material ID in ENVI-met before being integrated into the INX format. Table 9 in Appendix summarizes mapping of voxel classes in VoxCity to ENVI-met material and vegetation IDs in this study. Additionally, users can specify the trunk height ratio relative to total canopy height and the leaf area density (LAD) of trees. These tree-specific settings are then exported separately as a project database (EDB) file.
Figure 9 shows microclimate simulation examples using the widely-used ENVI-met code (V5.7.1, ENVI-met Gmbh, Essen) in two characteristic sites, high-rise Al Zahiya district in Abu Dhabi and low-rise residential in Greenbelt, MD (USA). Since ENVI-met is based on the structured mesh (i.e., Cartesian coordinates), VoxCity can provide basic 3D geometries (Figures 9a and 9b) for the simulation. The calculation domains were 800 m × 800 m × 310 m for Abu Dhabi, and 500 m × 500 m × 50 m for Greenbelt. They comprised inner areas with buildings and trees with horizontally homogeneous grid spacing (4 m for Abu Dhabi and 2 m for Greenbelt here), which is off-the-shelf from VoxCity. Because perimeter margins are necessary for ENVI-met simulation to avoid unphysical results (as detailed in their documentation: https://envi-met.com/tutorials-plugins-faqs-helpful-info/), VoxCity provides a function to remove perimeter buildings and trees. For vertical grid spacing, VoxCity provides options that use the same size for the horizontal grid or set an arbitrary number of layers. Here, we set 19 layers for the Greenbelt case and 50 for the Abu Dhabi case. One should note that before running the simulation, ENVI-met automatically resizes the meshes within the five layers near the ground to a finer resolution. We also note that ENVI-met V5.7.1 often predicts unphysical results for complex terrains (as detailed on their website: http://www.envi-hq.com/), so we strongly recommend VoxCity users not to use DEM when conducting ENVI-met simulation.
VoxCity automatically sets default values of the thermal properties of buildings, ground, and trees by providing the INX file. Although overwriting the INX file, e.g., setting detailed building surface materials, can give more accurate predictions, we show the results with the default values here to verify their adequacy. For meteorological forcings, we used the 2004–2018 typical meteorological year data of the EnergyPlus Weather file (https://www.ladybug.tools/epwmap/) and selected the nearest weather stations for each site.
Figures 9c–9f show thermal environments at 1.5 m height on typical summer afternoons (2:00 p.m.), namely the horizontal wind vectors and the universal thermal climate index (UTCI) (derived from the air temperature, humidity, radiation, and wind speed) predicted by ENVI-met. In the Abu Dhabi case, the inlet wind was the northwesterly sea breeze from the Persian Gulf. The speed was relatively high above the roads and the open areas and reached 5 m/s at the maximum (Figure 9c), probably because the city block pattern is along the direction of the inlet wind. However, the inlet air temperature was 44°C, hotter than the human body, and therefore, as seen around point ① in Figure 9e for example, the increase in wind speed rather increased UTCI by convective heat transfer (see Fig. 7 of Bröde et al. [155]). Shading effects of buildings and trees on UTCI are evident, and the difference between sunlit and shaded areas reached >> 20 °C (Figure 9e). Weak and complex diverged flows with lower air temperature and UTCI values are found in the leeward of the buildings (e.g., around building ② in Figures 9c and 9e and the southeast city block), indicating cooler above-roof air moved to the ground level by the downdrafts. In the Greenbelt case, the wind speed was generally lower than the inlet value (1.5 m/s) due to the momentum sink by trees (Figure 9d). The UTCI distribution seems to correspond simply to the shading pattern, and the contrast between roads and under-tree areas is evident (Figure 9f). We verified the validity of various parametric scenarios in Abu Dhabi, Greenbelt, and a few other sites. The calculations were successfully completed in all cases, while physically unrealistic results were obtained when, as mentioned above, OSM’s LULC data included water surfaces or when using DEM data. Note that the validation results and limitations of the recent ENVI-met code (version 4 or above) are explained in the literature [64, 156, 152].